OpenStudy (anonymous):

A compound contains 69.94 percent iron and 30.06 percent oxygen. What is its molecular formula if the molar mass of the compound is 199.55 grams per mole?

OpenStudy (anonymous):

The percentage mass becomes the mass of reaction. 1 step: find the number of moles using the formula n=m/M , where n is number of moles, m is mass of reaction and M is molar mass( tru mass of the element). Do this for each and every element. For Fe - 69.94:56=1.25 For O: 30.06/16= 1.88 2 step: find the smallest answer and use every answer from step 1 to divide the smallest number For Fe: 1.25/1.25= 1 For O: 1.88-1.25=1.504, approximately 2. So empirical formula is FeO2 Step 3- find the molar mass of the compound: 56+16*2= 88g 4step - find the molar mass ratio: 199.55/88, approximately 2. 5 step- multiply this number with you empirical formula, having 2(FeO2)= Fe2O4.